OpenStudy (anonymous):

Find all the zeroes of the equation. -3x4+27x2+1200=0 I need to learn how to do it! Just not the answer please

OpenStudy (anonymous):

finding the zeroes is the same as solving for X order of operations first multiplication -3*4=-12, 27*2=54 -12+54+1200=0 now add everything else together The answer would be 1242

myininaya (myininaya):

I think 4 and 2 are exponents on x

myininaya (myininaya):

\[-3x^4+27x^2+1200=0\]

myininaya (myininaya):

try replacing x^2 with u to transform the equation into: \[-3u^2+27u+1200=0\]

myininaya (myininaya):

this is a quadratic equation

myininaya (myininaya):

use the quadratic formula

myininaya (myininaya):

you could also simplify the equation before using the formula

myininaya (myininaya):

divide both sides by -3

OpenStudy (anonymous):

Yeah the 4 and 2 are exponents

OpenStudy (anonymous):

oh, oops, sorry >.<

OpenStudy (anonymous):

Nah you're good! thanks for the help though :)

OpenStudy (anonymous):

oh wait never mind, I got it.